# Vibe coding: programming through conversation with artificial intelligence

### v1

@@ 3.8.1 What is the nature of trust in vibe coding and how is it developed?
Trust in vibe coding appears to be granular, dynamic, contingent on vigilant review, and evolves through interaction with the system, manifesting as a balance between convenience and caution.
@@
YT21 presents a practitioner with what appears to be a well-calibrated trust mindset. This developer articulates that he can “quickly get 80-90% of the way there with AI”. However, this trust is contingent, as he explicitly states, “I like to review the changes because it helps me remain in control to some extent.” The risk of overtrust is actively mitigated as the developer frequently rejects AI suggestions that use non-existent APIs or approaches he considers less maintainable. Frequent (albeit lightweight) code review was observed across all participants. Thus, reviewing changes appears to be important for vibe coding, not only for maintaining understanding of the code, but also for maintaining agency, authorial ownership, and trust.
YT22 further reinforces this notion of contingent trust. The creator explicitly states that he doesn’t believe in “blindly following the AI”, stating the importance of his reading and reviewing the code. He further states: “AI is just a tool.” This invocation of the tool-user relationship can be viewed as being in explicit contrast to delegation of responsibility. The developer consistently reviews code, compares it to documentation, and tests frequently, showing that trust is continuously re-established rather than statically assumed.
TW1 illustrates the nature of trust in relation to expertise. Her expertise allowed her to recognise “way too complicated code” generated by the model and to fix it, while speculating that some vibe coders may not have the expertise to address the issue themselves as she did. This suggests that trust in vibe coding may be significantly influenced by the user’s own technical expertise. When discussing a complex codebase of “150 thousand lines of good code” the creator emphasises that she cannot allow “bad” code (e.g., overly complex, not performant) to get into the repository. Even on “simpler” features, such as an information page that describes the benefits of a VIP membership to the website, TW1 still notes she will review the pull request before committing and deploying to her website. This was then challenged by a viewer saying that “code review is not vibe coding”, which suggests a tension between trust and verification in the vibe coding paradigm.
@@ 3.9.1 How do programmers define vibe coding?
Vibe coding thus appears as a programming approach characterised by conversational, iterative interaction with AI tools, where the model handles significant portions of the coding work. From YT15, we see vibe coding defined through spoken, relatively casual yet detailed prompts and a process where the model is the primary agent for code creation and refinement. Similarly, in TW1, a programmer directly defines it as having “the AI do the heavy lifting” and describes the workflow as having the AI creating a feature and then “spruce it all up.” In YT21, the programmer describes vibe coding as “just chatting with the app, saying do this, do that”.
@@
The position on this spectrum also appears to vary based on individual preference and project phase. TW1’s comment about “don’t review just submit” (albeit humorously, as TW1’s workflow relied heavily on reviewing AI generated code) suggests some practitioners might push toward greater AI reliance, while others, like those in YT21 and YT22, maintain more balanced approaches. YT15 shows that programmers might be more willing to accept AI-generated code quickly in early stages but become more critical as development progresses. In YT22, a programmer initially describes vibe coding as when “you don’t really dig into the code, you just embrace the AI,” suggesting minimal human intervention. However, the same programmer later clarifies that he “doesn’t believe in blindly following the AI” and emphasises the importance of reading and reviewing code. This contradiction reflects the evolving and personal nature of how vibe coding is defined.
@@ 3.9.2 How does the performative aspect of vibe coding on streaming platforms influence the practice?
Creators frequently emphasise the impressive capabilities of AI through enthusiastic reactions, with YT15’s programmer repeatedly saying “wow”, such as when the AI-generated code for animations worked correctly. In YT21, the creator explicitly highlights that “because of the capabilities of AI and Cursor, [he] was able to accomplish [building the app],”. Creators emphasise the ease and speed of vibe coding (e.g. “you can develop this code even without touching a keyboard”, “boost my productivity”), invoke rhetorical flourishes like “trust the vibe” and call the results “awesome” or “incredible”. Even obvious failures may be framed positively (e.g., in one instance a creator hypothesised that transcription errors force the model to “think more broadly”, becoming a potential asset). These reactions amplify the perception of AI effectiveness for audiences, potentially to the point of exaggeration.
Streaming workflows introduces unique dynamics, as seen in TW1, where the creator frequently responds to chat questions and comments which can cause tangents commonly found in streaming workflows (Drosos and Guo,, 2021) including playing a game on her website because a viewer mentioned wanting an specific achievement. These interactive elements generate spontaneous content that wouldn’t occur in private coding but can yield valuable insights, as when she explained that “the context window is big now so [she doesn’t] have to worry about [token usage] as much” in response to a viewer’s concern about instruction length.
@@
Despite the emphasis on showcasing AI capabilities, the performances still demonstrate that realistic workflows include debugging and manual intervention. In YT22, while emphasizing “embracing the AI,” the creator openly shows instances of manually correcting AI output, updating Next.js versions, deleting unwanted features, and fixing typos. This provides a more authentic portrayal of the vibe coding process through acknowledging and detailing the necessary human oversight.
Moreover, creators often verbalise their understanding and expectations when they are somewhat uncertain. Thus, this running commentary might not just be purely directed at impression management (discussed later in Section 4) but also a way for the creator to maintain a sense of control and momentum during development. By narrating the process, they might be developing and iterating on their own mental model.
@@ 4.1 How vibe coding differs from previous generations of AI-assisted programming
The initial phase of AI-assisted programming tools, exemplified by early analyses of tools like GitHub Copilot, marked a significant shift in how programmers interact with code generation. Research from this period, such as the studies by Sarkar et al., 2022a , Vaithilingam et al., (2022) and Barke et al., (2023), documented how the programmer’s experience changes in response to AI capabilities like code completion based on context and comments. The subsequent generation of tools expanded upon these capabilities by evolving the scope and size of code generation within a codebase, and allowing the tools to access tools that expand the scope of their operations (e.g., running terminal commands to create files, start servers, install packages, etc.). This shift in capabilities is sometimes referred to as making these tools more “agentic”. Correspondingly, the experience of vibe coding, as observed in our videos, builds upon the experience of prior generations of AI-assisted programming, but introduces distinct nuances in philosophy, workflow, and the programmer’s engagement with the AI.
@@
A primary difference lies in the philosophy guiding the interaction. Vibe coding often embraces a more trust-based, hands-off approach, prioritizing flow or “vibe” over strict control. The explicit inspiration from the Karpathy canon encourages relying on the LLM to autonomously handle errors and work around difficult problems, moving away from traditional debugging methods. This introduces a workflow where the programmer may intentionally allow the AI to “drive” more consistently, even for familiar tasks, creating an interaction style that feels more conversational or like directing a capable assistant, distinct from the segmented modes previously identified (Barke et al.,, 2023).
@@
Prompting takes on a new character in vibe coding. Earlier research noted the difficulties users faced in effectively communicating intent and matching their abstraction level to the model’s capabilities, often relying on explicit comments (Sarkar et al., 2022a, ; Barke et al.,, 2023). Vibe coding appears to complexify this, featuring prompts that are often significantly longer and exhibit extreme mixing of granularity. Vibe coding prompts can blend vague aesthetic desires, broad ideological goals (like “make it 10X better”), specific technical constraints (like using a particular library or file location), and may even incorporate direct code snippets or external documentation links for context. Early tools primarily relied on comments or explicit pop-up input fields for prompting (Barke et al.,, 2023), while vibe coding integrates voice, text, and potentially other modalities in a fluid, less structured manner.
@@
The process of evaluation and debugging also shows differences. While earlier AI-assisted programming required validation through examination, execution, documentation lookup, and static analysis (Sarkar et al., 2022a, ; Barke et al.,, 2023), vibe coding prioritises rapid, targeted inspection. As detailed in Section 3.5.1, programmers quickly scan code diffs, looking for familiar patterns or keywords (eyeballing, glancing) rather than performing line-by-line reviews. Crucially, traditional manual debugging skills remain essential in vibe coding. When errors occur, vibe coders frequently revert to conventional methods: analysing error messages, using browser developer tools or the terminal, and forming their own hypotheses about the bug and its solution. The model might then be tasked with implementing the specific fix identified by the programmer, although sometimes raw error messages are simply provided to the model for repair. This workflow where the human diagnoses and plans, and the AI executes the fix, contrasts with the user experience of understanding and repairing AI code noted in earlier studies (Vaithilingam et al.,, 2022). It also differs from the explicit validation strategies like detailed examination and documentation lookup observed more prominently in the exploration mode of earlier tools (Barke et al.,, 2023).
@@
Vibe coding appears as an evolution of first-generation AI-assisted programming, that leans into the conversational and generative power of large language models while still requiring significant human expertise and judgment. It represents a move from code-generating models as an advanced autocomplete or search tool to a more integrated and capable development tool. Previous work had already identified the shortcomings of analogies between AI-assisted programming and programming via search-and-reuse, or compilation, or programming by specification (Sarkar et al., 2022a, ); these analogies are even weaker in the case of vibe coding.
@@ 4.2 A gestalt theory of vibe
In vibe coding, prompts are often high-level and describe the desired outcome rather than the specific implementation (shifting the focus of programmer intention and articulation towards the whole rather than the parts that compose it). The model then generates code, which, in conjunction with the running application and the context momentum that builds through iteration, can be seen as an emergent Gestalt.
@@
As detailed in Section 3.5.1, programmers spend less time deeply understanding every line generated by the model and more time validating its output against their high-level expectations and mental model. The ability to do this appears to be profoundly dependent on the programmer’s existing expertise and well-established mental schemas of coding patterns, frameworks, and their specific codebase. This expertise goes beyond the expertise required to write code manually. We posit that this is expertise in understanding the gestalt of the code and the context of the conversational session with the model. Programmers repeatedly exhibited the ability to look at a block of code and immediately identify its purpose within the overall program, and identify potential problems or inefficiencies quickly. This is possible because the code conforms to a mental schema that the programmer possesses, and is capable of rapidly updating as the code evolves, so that the vibe gestalt continues to be useful for future steering and verification. This expert schema, unlike in traditional programming, goes beyond the well-documented expert ability to read and synthesise a code base quickly, because it incorporates several novel constituent elements in the gestalt: the AI-generated commentary, the conversational chat history and context momentum, and knowledge and prior expectations about model capabilities.
@@ 4.3 The consequences of material disengagement from code
Nevertheless, the iteration central to vibe coding demonstrates how engagement with the model’s material responses acts as a form of resistance that actively shapes and refines the programmer’s intentions and strategies. The unexpected outputs, errors, or failures to meet requirements function as the resistances or affordances of this new mediating material, pushing back against the programmer’s initial intent and necessitating adjustments or refinements. This dynamic aligns with the concept of designing “productive resistances” (Sarkar,, 2024) in AI tools to cultivate intention, counteracting tendencies towards mechanised convergence or numbed intentionality. While vibe coding workflows often seek to minimise friction, the inherent pushback from the model’s outputs, and the programmer’s active evaluation and response to these, constitutes a form of material engagement that shapes intentions. Intentionally designed productive resistances could formalise and enhance a process already occurring organically within vibe coding workflows.
@@
4.3.1 Vibing beyond coding
As touched upon in Section 1, the vibe coding workflow, where the user avoids direct manipulation of the “raw” material substrate (e.g., text in a document, formulas in a spreadsheet, paint on a canvas) may spread to other AI-assisted knowledge workflows. However, the potential for turning these workflows into “vibe” versions of themselves may differ between domains. The programming context has resources that the practitioner can draw upon to support their understanding of the vibe gestalt that do not easily transfer to other domains, such as the saliency of identifiers and component structure. Code can be tested through unit tests or throw exceptions to aid detection of improperly generated code. Code can also be compiled and deployed for quick visual and interactive inspection by the user to judge whether the AI has been effective at code production. Text, on the other hand, is altogether more complex. Auditing generated text may inevitably require line-by-line reading to ensure that a vibe-written document is accurate and reflects the user’s intent. Therefore, expanding vibe coding workflows to other domains may require creating structural support for users to audit “vibey” AI-assisted workflows (Gordon et al.,, 2024) before we see vibing effectively expand to domains beyond coding.
@@ 4.4 Impression management and self-presentation: vibing for an audience
Reif et al., (2025) provides empirical evidence for a social evaluation penalty associated with using AI tools in the workplace. Their studies show that people anticipate and receive negative judgments from others, being perceived as lazier, less competent, and less diligent than those who use non-AI tools or no help at all. This penalty is linked to observers making negative dispositional inferences, attributing the use of AI to personal deficits rather than situational factors. There is thus a dilemma where the productivity benefits of AI can come at a social cost, leading some employees to conceal their AI use.

### v2

@@ 3.8.1 What is the nature of trust in vibe coding and how is it developed?
Trust in vibe coding appears to be granular, dynamic, contingent on review, and evolves through interaction with the system, manifesting as a tension between efficiency and comprehension.
@@
YT21 presents a practitioner with what appears to be a well-calibrated trust mindset. This developer articulates that they can “quickly get 80-90% of the way there with AI”. However, this trust is contingent, as they explicitly state, “I like to review the changes because it helps me remain in control to some extent.” The risk of overtrust is actively mitigated as the developer frequently rejects AI suggestions that use non-existent APIs or approaches they considers less maintainable. Frequent (albeit lightweight) code review was observed across all participants. Thus, reviewing changes appears to be important for vibe coding, not only for maintaining understanding of the code, but also for maintaining agency, authorial ownership, and trust.
YT22 further reinforces this notion of contingent trust. The creator explicitly states that they don’t believe in “blindly following the AI”, stating the importance of reading and reviewing the code. They further state: “AI is just a tool.” This invocation of the tool-user relationship can be viewed as being in explicit contrast to delegation of responsibility. The developer consistently reviews code, compares it to documentation, and tests frequently, showing that trust is continuously re-established rather than statically assumed.
TW1 illustrates the nature of trust in relation to expertise. Their expertise enables the recognition of “way too complicated code” generated by the model and enables them to fix it, while speculating that “the average vibe coder might not be able to fix that” without expertise such as TW1 possessed. This suggests that trust in vibe coding may be significantly influenced by the user’s own technical expertise. When discussing a complex codebase of “150 thousand lines of good code” the creator emphasises that they cannot allow “bad” code to enter the repository. Even on “simpler” features, such as an information page that describes the benefits of a VIP membership to the website, TW1 still notes they will review the pull request before committing and deploying to the website. Interestingly, this disposition was challenged by a viewer of TW1’s twich stream, saying that “code review is not vibe coding”, which suggests a tension between trust and verification in the vibe coding paradigm.
@@ 3.9.1 How do programmers define vibe coding?
Vibe coding thus appears to be a programming approach characterised by conversational, iterative interaction with AI tools, where the model handles significant portions of the coding work. TW1 directly defines it as having “the AI do the heavy lifting” and describes the workflow as having the AI creating a feature and then “spruce it all up.” In YT21, the programmer describes vibe coding as “just chatting with the app, saying do this, do that”.
@@
The position on this spectrum also appears to vary based on individual preference and project phase. TW1’s comment: “don’t review just submit” (albeit humorously, as TW1’s workflow relied heavily on reviewing AI-generated code), suggests some practitioners might push toward greater AI reliance, while others, like those in YT21 and YT22, maintain more balanced approaches. YT15 shows that programmers might be more willing to accept AI-generated code quickly in early stages but become more critical as development progresses. YT22 initially describes vibe coding as when “you don’t really dig into the code, you just embrace the AI,” suggesting minimal human intervention. However, the same programmer later clarifies that they “don’t believe in blindly following the AI” and emphasises the importance of reading and reviewing code. This contradiction reflects the evolving and personal nature of how vibe coding is defined.
@@ 3.9.2 How does the performative aspect of vibe coding on streaming platforms influence the practice?
Creators frequently emphasise the impressive capabilities of AI through enthusiastic reactions, with YT15’s programmer repeatedly saying “wow”, such as when the AI-generated code for animations worked correctly. In YT21, the creator explicitly highlights that “because of the capabilities of AI and Cursor, [the creator] was able to accomplish [building the app],”. Creators emphasise the ease and speed of vibe coding (e.g. “you can develop this code even without touching a keyboard”, “boost my productivity”), invoke rhetorical flourishes like “trust the vibe” and call the results “awesome” or “incredible”. Even obvious failures may be framed positively (e.g., in one instance a creator hypothesised that transcription errors force the model to “think more broadly”, becoming a potential asset). These reactions amplify the perception of AI effectiveness for audiences, potentially to the point of exaggeration.
Streaming workflows introduces unique dynamics, as seen in TW1, where the creator frequently responds to chat questions and comments which can cause tangents commonly found in streaming workflows (Drosos and Guo,, 2021) including playing a game on their website because a viewer mentioned wanting an specific achievement. These interactive elements generate spontaneous content that wouldn’t occur in private coding but can yield valuable insights, as when they explained that “the context window is big now so [they don’t] have to worry about [token usage] as much” in response to a viewer’s concern about instruction length.
@@
Despite the emphasis on showcasing AI capabilities, the performances still demonstrate that realistic workflows include debugging and manual intervention (as elaborated in Sections 3.5 and 3.7.3). Moreover, creators often verbalise their understanding and expectations when they are uncertain. Thus, this running commentary might not just be purely directed at impression management (discussed later in Section 4) but also a way for the creator to maintain a sense of control and momentum during development. By narrating the process, they might be developing and iterating on their own mental model.
@@ 4.1 How vibe coding differs from previous generations of AI-assisted programming
The initial phase of AI-assisted programming tools, exemplified by early analyses of tools like GitHub Copilot, marked a significant shift in how programmers interact with code generation. Research from this period, such as the studies by Sarkar et al., 2022a , Vaithilingam et al., (2022) and Barke et al., (2023), documented how the programmer’s experience changes in response to AI capabilities like code completion based on context and comments. The subsequent generation of tools expanded upon these capabilities by evolving the scope and size of code generation within a codebase, and allowing the language model to access other tools that expand the scope of their operations (e.g., running terminal commands to create files, start servers, install packages, etc.). This shift in capabilities is sometimes referred to as making these tools more “agentic”. Correspondingly, the experience of vibe coding, as observed in our videos, builds upon the experience of prior generations of AI-assisted programming, but introduces distinct nuances in philosophy, workflow, and the programmer’s engagement with the AI.
@@
A primary difference lies in the philosophy guiding the interaction. Vibe coding often embraces a more trusting, hands-off approach, prioritising flow or “vibe” over strict control. The explicit inspiration from the Karpathy canon encourages relying on the LLM to autonomously handle errors and work around difficult problems, moving away from traditional debugging methods. This introduces a workflow where the programmer may intentionally allow the AI to “drive” more consistently, even for familiar tasks, creating an fluid interaction style that is distinct from the segmented modes previously identified (Barke et al.,, 2023).
@@
Prompting takes on a new character in vibe coding. Earlier research noted the difficulties users faced in effectively communicating intent and matching their abstraction level to the model’s capabilities, often relying on explicit comments (Sarkar et al., 2022a, ; Barke et al.,, 2023). Vibe coding appears to complicate this, featuring prompts that are often significantly longer and exhibit extreme mixing of granularity. Vibe coding prompts can blend vague aesthetic desires, broad ideological goals (like “make it 10X better”), specific technical constraints (like using a particular library or file location), and may even incorporate direct code snippets or external documentation links for context. Early tools primarily relied on comments or explicit pop-up input fields for prompting (Barke et al.,, 2023), while vibe coding integrates voice, text, and potentially other modalities in a fluid, less structured manner.
@@
The process of evaluation and debugging also shows differences. While earlier AI-assisted programming required validation through examination, execution, documentation lookup, and static analysis (Sarkar et al., 2022a, ; Barke et al.,, 2023), vibe coding prioritises rapid, targeted inspection. As detailed in Section 3.5.1, programmers quickly scan code diffs, looking for familiar patterns or keywords (eyeballing, glancing) rather than performing line-by-line reviews. Crucially, traditional manual debugging skills remain essential in vibe coding. When errors occur, vibe coders frequently revert to conventional methods: analysing error messages, using browser developer tools or the terminal, and forming their own hypotheses about the bug and its solution. The model might then be tasked with implementing the specific fix identified by the programmer, although sometimes raw error messages are simply provided to the model for repair. This workflow where the human diagnoses and plans, and AI executes the fix, contrasts with the user experience of understanding and repairing AI code noted in earlier studies (Vaithilingam et al.,, 2022). It also differs from the explicit validation strategies like detailed examination and documentation lookup observed more prominently in the exploration mode of earlier tools (Barke et al.,, 2023).
@@
Vibe coding is thus an evolution of first-generation AI-assisted programming, that leans into the conversational and generative power of large language models, while still requiring significant human expertise and judgment. It represents a move from code-generating models as an advanced autocomplete or search tool to a more integrated and capable development tool. Previous work had already identified the shortcomings of analogies between AI-assisted programming and programming via search-and-reuse, or compilation, or programming by specification (Sarkar et al., 2022a, ); these analogies are even weaker in the case of vibe coding.
@@ 4.2 A gestalt theory of vibe
In vibe coding, prompts are often high-level and describe the desired outcome rather than the specific implementation (shifting the focus of programmer intention and articulation towards the whole rather than the parts that compose it). The model then generates code, which, in conjunction with the running application and the context momentum (Section 3.2.2) that builds through iteration, can be seen as an emergent gestalt.
@@
As detailed in Section 3.5.1, programmers spend less time deeply understanding every line generated by the model and more time validating its output against their high-level expectations and mental model. The ability to do this appears to be profoundly dependent on the programmer’s existing expertise and well-established mental schemas of coding patterns, frameworks, and their specific codebase. This expertise goes beyond the expertise required to write code manually. We posit that this is expertise in understanding the gestalt of the code within the context of the conversational session with the model. Programmers repeatedly exhibited the ability to look at a block of code and immediately identify its purpose within the overall program, and identify potential problems or inefficiencies quickly. This is possible because the code conforms to a mental schema that the programmer possesses, and is capable of rapidly updating as the code evolves, so that the vibe gestalt continues to be useful for future steering and verification. This expert schema, unlike in traditional programming, goes beyond the well-documented expert ability to read and synthesise a code base quickly, because it incorporates several novel constituent elements in the gestalt: the AI-generated commentary, the conversational chat history and context momentum, and knowledge and prior expectations about model capabilities.
@@ 4.3 The consequences of material disengagement from code
Nevertheless, the iteration central to vibe coding demonstrates how engagement with the model’s material responses acts as a form of resistance that actively shapes and refines the programmer’s intentions and strategies. The unexpected outputs, errors, or failures to meet requirements function as the resistances or affordances of this new mediating material, pushing back against the programmer’s initial intent and necessitating adjustments or refinements. While vibe coding workflows often seek to minimise friction, the inherent pushback from the model’s outputs, and the programmer’s active evaluation and response to these, constitutes a form of material engagement that shapes intentions.
This dynamic aligns with the concept of designing “productive resistances” (Sarkar,, 2024) in AI tools to cultivate intention, counteracting tendencies towards weakened intentionality, and thereby mechanised convergence (Sarkar, 2023a, ) of outcomes. Intentionally designed productive resistances could formalise and enhance a process already occurring organically within vibe coding workflows.
@@
Vibing beyond coding.
As touched upon in Section 1, the vibe coding workflow, where the user avoids direct manipulation of the “raw” material substrate (e.g., text in a document, formulas in a spreadsheet, paint on a canvas) may spread to other AI-assisted knowledge workflows. However, the potential for turning these workflows into “vibe” versions of themselves may differ between domains. The programming context has resources that the practitioner can draw upon to support their understanding of the vibe gestalt that do not easily transfer to other domains, such as the saliency of identifiers and component structure. Code can be tested through unit tests or throw formal exceptions to aid detection of errors. Code can also be compiled and deployed for visual and interactive inspection by the user to judge whether code generation has been satisfactory.
Text, on the other hand, is altogether more complex. Auditing generated text may inevitably require line-by-line reading to ensure that a “vibe written” document is accurate and reflects the user’s intent. Therefore, expanding vibe coding workflows to other domains may require creating structural support for users to co-audit (Gordon et al.,, 2024) AI-intermediated workflows before we see vibing effectively expand to domains beyond coding.
@@ 4.4 Impression management and self-presentation: vibing for an audience
To understand how our participants engage in impression management, it is necessary to contextualise our observations within the emerging understanding of the social dynamics of AI use. Reif et al., (2025) provides empirical evidence for a social evaluation penalty associated with using AI tools in the workplace. Their studies show that people anticipate and receive negative judgments from others, being perceived as lazier, less competent, and less diligent than those who use non-AI tools or no help at all. This penalty is linked to observers making negative dispositional inferences, attributing the use of AI to personal deficits rather than situational factors. There is thus a dilemma where the productivity benefits of AI can come at a social cost, leading some employees to conceal their AI use.